OpenStudy (anonymous):

Use a calculator to approximate 2 values of feta (positives only) of cos[feta]=.8746 Round to 2 decimal places.

OpenStudy (agent0smith):

Do you mean theta? cos(feta) = cheese (...because feta = cheese) \[\cos \theta = 0.8746\] take the inverse cosine of both sides \[\cos^{-1} \cos \theta = \cos^{-1} 0.8746\] \[\theta = \cos^{-1} 0.8746\]

OpenStudy (anonymous):

Ahaha yes sorry. I did that and I got 29 degrees and from previous problems you're suppose to add or subtract 180 since they asked for 2 values. I'm confused on that part.

OpenStudy (agent0smith):

Since you want positives, you can add multiples of 360, or subtract 29 degrees from 360.

OpenStudy (anonymous):

Wait, I'm having a hard time understanding..why not 180?

OpenStudy (agent0smith):

Try adding 180 to 29... cos209 = -0.8746 (not 0.8746)

OpenStudy (anonymous):

OHHHHHHHHH THANK YOU SO MUCH!!

OpenStudy (agent0smith):

No prob. cos (360° − x) = cos x but cos (180° + x) = − cos x

OpenStudy (anonymous):

Thanks!! That^ helped alot too.

OpenStudy (agent0smith):

and also cos (360° + x) = cos x That one applies to sine/tan too (since adding multiples of 360 is just going around the circle in whole rotations, so the 'reference angle' stays the same)

OpenStudy (anonymous):

Thanks :)